OpenStudy (anonymous):

evaluate the expression without using a calculator log 7 1/7

OpenStudy (anonymous):

this is saying how is 7=1/7?

OpenStudy (anonymous):

yu need to fins the power at which 7^n=1/7. What can be n?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

\[k ^{-1}= \frac{ 1 }{ k }\]

OpenStudy (anonymous):

yes. So n = -1. That's is the answer